Wendy and Charlotte both enjoy watching Manu’s Kitchen Rules (or MKR). Initially, assume that when broadcast on public television MKR is a public good that is both non-excludable and non-rivalrous. Wendy and Charlotte have demand for MKR given by the following:

Qw=122 − 2

0 0
Answer #1

Market demand Q=162-3p

Marginal cost mc=50

Under competitive economy, p=mc=50

Quantity produced q= 162- 3(50)= 12

Efficient level of production is 12

Qw= 122-2(50)= 22

Qc= 40-50=-10

Total consumption Qw+Qc= 12

b) if Mc= 24

Under competitive economy, p=mc=24

Qw= 122-2(24)= 74

Qc= 40- 24= 16

Total produced Qw+Qc= 90 units consumed in the competitive equilibrium.

c) third degree price discrimination leads to profit maximization from individual demand function

Marginal revenue of wendy(mrw)= marginal cost

mrw=122-4p=mc=24

pw=29.5 price charged from Wendy

Marginal revenue from Charlotte(mrc)=marginal cost

mrc= 40-2p=mc = 24

pc=8 price charged from Charlotte